You buy 6 bags of bird food. sunflower seeds sells for $5.oo a bag. a bird feed mix is $4.00 a bad. if you spent $28.00, how many bags of each type did you buy?
Answer by fractalier(6550)

Call the number of bags of seeds and bird feed, x and y. Then we write
x + y = 6 and
5x + 4y = 28
Now multiply the top equation by 5 and subtract it...we get
5x + 4y = 28
-(5x + 5y = 30)
---------------
-y = -2
y = 2 bags of bird feed mix
x = 4 bags of seeds